# CoLA: Exploiting Compositional Structure for Automatic and Efficient Numerical Linear Algebra

### v1

@@ 3.4 Automatic Differentiation and Machine Learning Readiness
Multi framework support and GPU/TPU acceleration  CoLA is compatible with both PyTorch and JAX. This compatibility not only makes our framework plug-and-play with existing implemented models, but it also adds GPU/TPU support, differentiating it from existing solutions (see Table 2).
@@ 4 Applications
Gaussian processes training  GPs are flexible nonparametric probabilistic models where inductive biases are expressed through a covariance (kernel) function. At its core, training a GP involves computing and taking gradients of the log determinant of a kernel $\log\left|{\bm{\mathbf{K}}}\right|$ and of a quadratic term ${\bm{\mathbf{y}}}^{T}{\bm{\mathbf{K}}}^{-1}{\bm{\mathbf{y}}}$ (where $\bm{\mathbf{y}}$} is the vector of observations).
@@ C.2 Iterative methods
${\bm{\mathbf{A}}}{\bm{\mathbf{x}}}={\bm{\mathbf{b}}}$ (non-symmetric) GMRES BiCGSTAB, CR, QMR
@@
$\texttt{Eigs}({\bm{\mathbf{A}}})$ (non-symmetric) Arnoldi IRAM, QR algorithm
$\texttt{Eigs}({\bm{\mathbf{A}}})$ (self-adjoint) Lanczos LOBPCG, Rayleigh-Ritz, Bi-Lanczos
$\bm{\mathbf{A}}$$^{+}$ CG CGS, LSQR, LGMRES
@@
$f({\bm{\mathbf{A}}})$ (self-adjoint) SLQ SVD, Rational Krylov Subspaces
Table 5: CoLA’s base case iterative algorithm and some alternatives. We now expand on the acronyms. GMRES: Generalized Minimum RESidual, BiCGSTAB: BiConjugate Gradient STABilized, CR: Conjugate Residuals, QMR: Quasi-Minimal Residual, MINRES: MINimum RESidual, CG: Conjugate Gradients, IRAM: Implicitly Restarted Arnoldi Method, LOBPCG: Locally Optimal Block Preconditioned Conjugate Gradients, Bi-Lanczos: Bidiagonal Lanczos, CGS: Conjugate Gradient Squared, LSQR: Least-Squares QR, LGMRES: Least-sqaures Generalized Minimum RESidual, SVD: Singular Value Decomposition, rSVD: randomized Singular Value Decomposition, and SLQ: Stochastic Lanczos Quadrature.
@@ D.3 Sum structure experiments
D.4 Applications

### v2

@@ 3.4 Automatic Differentiation and Machine Learning Readiness
Multi framework support and GPU/TPU acceleration  CoLA is compatible with both PyTorch and JAX. This compatibility not only makes our framework plug-and-play with existing implemented models, but it also adds GPU/TPU support, differentiating it from existing solutions (see Table 2). CoLA’s iterative algorithms are the class of linear algebra algorithms that benefit most from hardware accelerators as the main bottleneck of these algorithms are the MVMs executed at each iteration, which can easily be parallelized on hardware such as GPUs. Figure 3 empirically shows the additional impact of hardware accelerators across different datasets and linear algebra operations.
Figure 3: For sufficiently large problems, switching from dense to iterative algorithms provides consistent runtime reductions, especially on a GPU, where matrix multiplies can be effectively parallelized. We plot the ratio between the runtime of a linear algebra operation using CoLA or PyTorch on different hardware (CPU and GPU) divided by the runtime of using PyTorch CPU. For the linear solves, we use the matrix market sparse operator Trefethen; for the eigenvalue estimation, we use the matrix market sparse operator mhd4800b and, finally, for the log determinant computation, we use the matrix market sparse operator bcsstk18. We provide additional details in Section D.4.
@@ 4 Applications
Gaussian processes  GPs are flexible nonparametric probabilistic models where inductive biases are expressed through a covariance (kernel) function. At its core, training a GP involves computing and taking gradients of the log determinant of a kernel $\log\left|{\bm{\mathbf{K}}}\right|$ and of a quadratic term ${\bm{\mathbf{y}}}^{T}{\bm{\mathbf{K}}}^{-1}{\bm{\mathbf{y}}}$ (where $\bm{\mathbf{y}}$} is the vector of observations).
@@ 5 Discussion
Acknowledgements
This work is supported by NSF Award 1922658, NSF CAREER IIS-2145492, BigHat Biosciences, Capital One, and an Amazon Research Award.
@@ C.2 Iterative methods
${\bm{\mathbf{A}}}{\bm{\mathbf{x}}}={\bm{\mathbf{b}}}$ (non-symmetric) GMRES BiCGSTAB, LGMRES, QMR
@@
$\texttt{Eigs}({\bm{\mathbf{A}}})$ (non-symmetric) Arnoldi IRAM, Bi-Lanczos
$\texttt{Eigs}({\bm{\mathbf{A}}})$ (self-adjoint) Lanczos LOBPCG
$\bm{\mathbf{A}}$$^{+}$ CG LSQR, LSMR
@@
$f({\bm{\mathbf{A}}})$ (self-adjoint) SLQ Arnoldi
Table 5: CoLA’s base case iterative algorithm and some alternatives. We now expand on the acronyms. GMRES: Generalized Minimum RESidual, BiCGSTAB: BiConjugate Gradient STABilized, QMR: Quasi-Minimal Residual, MINRES: MINimum RESidual, CG: Conjugate Gradients, IRAM: Implicitly Restarted Arnoldi Method, LOBPCG: Locally Optimal Block Preconditioned Conjugate Gradients, Bi-Lanczos: Bidiagonal Lanczos, CGS: Conjugate Gradient Squared, LSQR: Least squares QR, LSMR: Least squares Minimal Residual iteration, LGMRES: Least squares Generalized Minimum RESidual, rSVD: randomized Singular Value Decomposition, and SLQ: Stochastic Lanczos Quadrature.
@@ D.3 Sum structure experiments
D.4 Hardware speed-up comparisons
This section pertains to the experiments of Figure 3. For all these experiments we computed the runtime reduction as a fraction between the time that it takes CoLA to run some linear algebra operation and PyTorch using the same hardware. As an example, assume that PyTorch takes 200 seconds to compute a solve using a CPU and 100 seconds to compute the same solve but now using a GPU. Moreover, assume that CoLA’s iterative algorithm takes 100 seconds to compute the same solve on a CPU and 40 seconds on a GPU. Thus, the runtime reduction would be $100/200=0.5\%$ for the CPU column whereas $40/100=0.4$ for the GPU column.
  1. 1.
    Solves. In this experiment we calculated the % runtime reduction when running torch.linalg.solve on the Trefethen $N=20K$ matrix market sparse operator. In this experiment, CG was run with a tolerance of $10^{-11}$ and a maximum number of iterations equal to the operator size.
  2. 2.
    Eigenvalue estimation. In this experiment we calculated the % runtime reduction when running torch.linalg.eigh on the mhd4800b $N=4.8K$ matrix market sparse operator. In this experiment, Lanczos was run with a tolerance of $10^{-9}$ and a maximum number of iterations equal to $100$.
  3. 3.
    Log determinant computation. In this experiment we calculated the % runtime reduction when running torch.linalg.logdet on the bcsstk18 $N=11.9K$ matrix market sparse operator. In this experiment, the stochastic Lanczos quadrature was run using $30$ Lanczos probe estimates and $25$ samples.
D.5 Applications